The table contains the Sales estimates for the next year.

Month Sales $ Month Sales $
Jan         19,034 July     99,711
Feb         5,265 Aug     80,177
Mar       80,177 Sep     19,034
Apr     99,711 Oct     80,177
May         19,034 Nov         5,265
June     5,265 Dec     99,711

Sales are collected as follows:
In the month of Sales: 43%
In the next month: 33%
After 2 months: 9%
After 3 months: remainder is collected

What is the cash collection for July?

Enter your answer rounded off to two decimal points. Do not enter $ or comma in the answer box.
Answer #1

Cash collections for July will be as follows:

July Sales = 99,711*43% = $42,875.73

June Sales = 5265*33% = $1737.45

May Sales = 19034*9% = $1713.06

Apr Sales = 99,711*15% = $14,956.65

Hence, total expected collections in July = $61,282.89